can someone show me how do this problem
Fahrenheit temperature. Water boils at 212°F.
a) Use the accompanying graph to determine the Celsius
temperature at which water boils.
b) Find the Fahrenheit temperature of hot tap water at
70°C by solving the equation 70=5/9(f-32)

Answer by Alan3354(69443)

can someone show me how do this problem
Fahrenheit temperature. Water boils at 212°F.
a) Use the accompanying graph to determine the Celsius
temperature at which water boils.
I don't see a graph, but 212º F is 100º C.
------------------------------
b) Find the Fahrenheit temperature of hot tap water at
70°C by solving the equation 70=5/9(f-32)
70 C = 1.8*70 + 32 = 158º F